Create a pseudocode and flowchart for the following:

  1. Convert Fahrenheit degree to Celsius
  2. Converting feet into meters
  3. Checking if integer 1 is equal to integer 2
  4. Check if student's grade is PASSED or FAILED. PASSED if grade is greater than or equal to 75. FAILED if grade is less than 75.
  5. Compute for the perimeter of a circle.
1
Expert's answer
2021-11-15T12:36:28-0500

input:

tf \\ Fahrenheit degree

tc \\ Celsius degree

ft \\ feet

m \\ meters

int a \\ integer 1

int b \\ integer 2

m1 \\ message of checking if integer 1 is equal to integer 2

g \\ student's grade

m2 \\ message of checking of student's grade

r \\ radius of a circle

p \\  perimeter of a circle


tc = (tf-32)*5/9 \\ Convert Fahrenheit degree to Celsius

m = ft/3.28 \\ Converting feet into meters


if a == b then \\ checking if integer 1 is equal to integer 2

m1 = 'integer 1 is equal to integer 2'

else

m1 = 'integer 1 is not equal to integer 2'


if g => 75 then \\ checking of student's grade

m2 = 'student's grade is PASSED'

else

m2 = 'student's grade is FAILED'


p = "2\pi r" \\ Compute for the perimeter of a circle


output:

tf 'Fahrenheit degree is' tc 'Celsius degree'

ft 'feet is' m 'meters'

m1

m2

'perimeter of a circle is' p